# Resource Estimation for Delayed Choice Quantum Entanglement Based Sneakernet Networks Using Neutral Atom qLDPC Memories

### v1

@@ Abstract
Quantum Entanglement is a vital phenomenon required for realizing secure quantum networks, so much that distributed entanglement can be re-imagined as a commodity which can be traded to enable and maintain these networks. We explore the idea of commercializing entanglement-based cryptography and future applications where advanced quantum memory systems support less advanced users. We design a sneakernet-based quantum communication network with a central party connecting the users through delayed-choice quantum entanglement swapping, using quantum Low-Density-Parity-Check (qLDPC) encoded qubits on neutral atoms. Our analysis compares this approach with traditional surface codes, demonstrating that qLDPC codes offer superior scaling in terms of resource efficiency and logical qubit count. We show that with near-term attainable patch sizes, one can attain medium-to-high fidelity correlations, paving the way towards large-scale commercial quantum networks.
@@ I Introduction
The second Quantum Revolution refers to a future technological period when all the features of quantum mechanics are harnessed for information processing devices, expanding beyond the limited application of quantum mechanics in current computers, communication systems, and sensors Dowling and Milburn (2003); Deutsch (2020); Jaeger (2018); Atzori and Sessoli (2019); Rohde (2021a). The ultimate achievement of this revolution is embodied by quantum computers that are capable of operating on a vast scale, with the ability to tolerate faults and fix errors. These computers would need millions or more logical qubits Riel (2021) to tackle complex cryptographic tasks. While the creation of such devices remains in its early stages, there are compelling grounds to anticipate that scalable quantum computers will become tangible at some point within this century Gambetta (2020).
@@
While there is reason to be hopeful about the progress of large-scale quantum computers, it remains uncertain if quantum computers for everyday consumers, such as a quantum smartphone, will ever become feasible. This refers to a compact and powerful computational device that is connected to a quantum network, possibly on a global scale. Possible constraints that could prevent the development of quantum smartphones include the requirement for extremely low temperatures to cool down qubits Gardiner and Zoller (2015), or the necessity for optical communication channels O’brien (2007) instead of microwave ones to connect to the quantum internet, the vulnerability of qubits to mechanical vibrations and the impact of the Earth’s magnetic field Körber et al. (2018). Suppose we accept the possibility of large-scale quantum computers and a quantum internet. In that case, it is logical to enquire how a customer, who has no access to a quantum memory, but only has access to conventional gear such as classical detectors, classical memory, and classical computing, can participate in and perform complex quantum protocols.
@@
One way is to do this by the use of distributed quantum entanglement Perseguers et al. (2013). Quantum entanglement is a pivotal phenomenon that forms the basis of various protocols in quantum communications and quantum computing Rohde (2021b); Nielsen and Chuang (2001); Horodecki et al. (2009); Wehner et al. (2018); Kimble (2008); Cacciapuoti et al. (2019). Due to it being a major primitive for most quantum tasks, it would be natural to consider quantum entanglement as a tradeable commodity, such that users can purchase it to establish connections with other users in the network to perform quantum tasks such as quantum key distribution, distributed quantum computing, quantum authentication, etc. Commoditizing entanglement is an essential step toward large-scale commercial quantum networks.
@@
To better demonstrate the above-mentioned ideas of entanglement commercialization and quantum access to common users with classical hardware, we design and analyse an entanglement one-time-pad QKD Sharma and Kumar (2019) based quantum network model networked using delayed-choice entanglement swapping Peres (2000). In this study, we demonstrate that delayed-choice entanglement swapping enables customers, who only own classical hardware, to use the offerings of a hypothetical future quantum memory company for the purpose of generating secure quantum keys. These keys are based on the E91 protocol Ekert (1991). Our design employs a central party Charlie, who generates Bell-pairs, loads the halves separately onto different qLDPC codes, and distributes one set of the halves via sneakernet Devitt et al. (2016); Zhong et al. (2015) to multiple endpoints where it is immediately measured upon arrival while keeping the other set of halves to himself on which he later performs Bell measurements to generate correlations between the endpoint measurements of any pair of users. This enables the users to execute this task in a versatile manner by pre-purchasing their key bits prior to selecting their communication partners. This protocol is beneficial for creating secure connections between networks of devices that have minimal hardware needs but can interact with a single, advanced quantum node or several nodes.
@@
We design our protocol on quantum Hyper Graph Product (HGP) codes Tillich and Zémor (2013), which are a class of Low-Density Parity-Check (qLDPC) codes Breuckmann and Eberhardt (2021) for quantum error correction Lidar and Brun (2013), implemented on a neutral atom architecture Henriet et al. (2020); Xu et al. (2024). These codes offer significant advantages over traditional surface codes Fowler et al. (2012) in terms of encoding rate and resource efficiency Bravyi et al. (2024); Xu et al. (2024). Our analysis provides a detailed comparison between qLDPC and surface codes, demonstrating that qLDPC codes allow us to achieve higher logical qubit counts with fewer physical qubits as they have constant encoding rates regardless of patch size. This property potentially reduces the hardware requirements and improves the scalability of our proposed system. We have selected neutral atoms as our physical platform, as their ability to shuttle and rearrange within the lattice enables selective qubit unloading and efficient syndrome extraction Xu et al. (2024). This capability also allows us to use only one set of ancilla qubits and a single surface code patch, positioned adjacent to the qLDPC code block, for both loading and unloading operations via teleportation Xu et al. (2024).
@@
The foundation of our study is based on the research conducted on consumer-level quantum key distribution Zhang et al. (2018); Lowndes et al. (2021); Duligall et al. (2006), as well as the research on quantum sneakernet Devitt et al. (2016), which explores the spread of entanglement across long distances with high bandwidth and low latency. While our primary focus is on secure communication protocols, it is feasible to adapt our methodology for many applications such as quantum voting Vaccaro et al. (2007), clock synchronization Giovannetti et al. (2001), and reference frame alignment Bagan et al. (2004).
@@ II Delayed Choice Entanglement Swapping
To describe our protocol, we start by providing a concise overview of the initial delayed-choice entanglement-swapping technique, as explained by Peres Peres (2000). In accordance with Peres, we examine the most basic iteration of this protocol. This protocol aims to establish entanglement and eventually generate correlated bits between two parties, Alice and Bob, who choose a third party, Charlie, to mediate the entanglement generation. Consider three individuals, namely Alice, Bob, and Charlie, as seen in the circuit diagram in Figure 1.
Figure 1: Circuit diagram of quantum entanglement swapping. In (a), the conventional entanglement swapping Coecke (2014) occurs, with Charlie performing the Bell measurement before Alice and Bob does. In (b), the order of measurement is reversed, with Alice and Bob measuring their qubits before Charlie. These two circuits are equivalent as the two sets of measurements are parallel, non-overlapping, and independent of each other.
Two Bell Pairs $q_{1}q_{2}$ and $q_{3}q_{4}$ are present such that $q_{1}$ is in the custody of Alice, $q_{4}$ with Bob, and $q_{2}$ and $q_{3}$ with Charlie. Qubits $q_{1}$ of Alice and $q_{4}$ of Bob get entangled when Charlie performs a joint Bell-measurement on his qubits $q_{2}$ and $q_{3}$. However, the specific relationship between Alice and Bob’s qubits cannot be known until Charlie reveals the outcome of his Bell state measurement. Alice and Bob’s photons are perfectly entangled, disregarding any errors or decoherence. This exact entanglement allows them to be used for various entanglement-based protocols, such as cryptography using the E91 protocol Ekert (1991).
@@
Peres’ idea goes beyond entanglement swapping by recognizing that Alice, Bob, and Charlie might potentially be space-like separated. This suggests that one can discover a frame of reference when the sequence of qubit measurement is modified, as shown by the Gisin group Zbinden et al. (2001). In the preceding discourse, we used the assumption that Charlie conducted his measurement before Alice and Bob did. Alternatively, it is possible for Alice and Bob to conduct their measurements prior to Charlie (Figure 1(b)). In this scenario, the outcome of their measurements itself is not affected by whether or not Charlie performed either the Bell state or independent measurement, but the degree of correlation between Alice and Bob’s measured bits is dependent upon the kind of measurement performed by Charlie. This serves as the foundation for the delayed-choice entanglement-swapping protocol and is crucial for the advancement of our ideas. In the delayed-choice entanglement swapping technique, Alice and Bob perform measurements on their qubits and record the outcomes using classical storage. At a later point, Charlie has a decision to either measure his qubits separately or do a Bell state measurement. The measurement outputs are either associated in a known manner or wholly uncorrelated, depending on the measurement option chosen by Charlie and the random basis choice of Alice and Bob.
@@ III Setup
Figure 2: (Inset) A schematic of a 5-node delayed choice quantum network with Charlie Inc. in the center. (Main) Delayed choice entanglement swapping protocol between two nodes Alice and Bob. S represents the surface code patches, A refers to the ancilla patches, C denotes Charlie’s qLDPC memory and T represents the qLDPC memory transported to the end-nodes.
@@
and is also described concisely in pseudocode 1. Charlie Inc., acting as a central node, first prepares Bell pairs in surface codes and teleports each half into separate qLDPC code blocks. This is done because direct computations and measurements for qLDPC codes are not yet well defined at the time of this writing. Charlie Inc. then keeps one set of neutral-atom encoded qLDPC blocks and sends the other set to an end-user via physical transport, like a train or a truck. This process is done for several such end-users at different destinations. This is called the ”sneakernet” method. At the user’s end, the user purchases a certain number of bits based on their needs. At the user’s end, a quantum ATM (qATM) “unloads” the incoming qLDPC blocks by sequentially teleporting the constituent qubits onto a surface code, followed by measuring the surface code randomly in X or Z bases, and commercially selling the resulting classical bits to users, which is in turn stored in a device belonging to the user, such as a smartphone. For our analysis, we assume that the process of transferring classical bits onto the end-users’ classical devices is secure. The measurement bases are then made public.
@@
Among several end-users in the network, let’s say two of them, namely Alice and Bob, want to connect at a later time. They decide upon how many bits they need to correlate for their requirement. Charlie Inc. then converts the corresponding logical qubits in the qLDPC blocks on his end into surface code patches and performs Bell measurements on those patches, thereby correlating Alice and Bob’s bits. Finally, Alice and Bob use their correlated bits to generate a draft encryption key. Depending on their security requirement, they can enhance the key’s security using CHSH tests Clauser et al. (1969) and privacy amplification, sacrificing some bits in the process. The remaining bits form a secure encryption key for end-to-end message encryption. This approach turns entanglement into a tradeable commodity and allows users with only classical hardware to participate in quantum cryptography protocols. For example, Alice measures a number of qubits, labeled ${A_{1},A_{2},...,A_{n}}$ and informs Charlie Inc. which qubit measurement results she now owns. Charlie Inc. then records the owner of the qubits and stores the other half of the Bell pairs in long-term quantum storage. Similarly, Bob buys the rights to measure qubits ${B_{1},B_{2},...,B_{n}}$ and Dave buys the rights to measure qubits ${D1,D_{2},...,D_{n}}$. At this stage, neither participant has decided who they wish to exchange correlations with, they have simply ‘topped up’ their future quantum potential i.e. the bits with which they can perform an unconditionally secure one-time-pad with another user of their choice in the future. If Alice and Dave wish to generate entangled data for some protocol, then they need to publicly declare to Charlie Inc. how many topped up bits they want to have correlated, then Charlie Inc. performs the entanglement swap between the appropriate (stored) Bell pairs and broadcasts the measurement results. If Alice and Dave wish to check Charlie Inc.’s fidelity, they may sacrifice some of their bits, and assuming they are satisfied, can proceed with standard key distillation and privacy amplification Bennett et al. (1995). If Alice then wishes to share secrets with Bob, then again, she need only declare to Charlie Inc. how many bits she wants to correlate (of course, she cannot use measurement results that have previously been used with Dave), and proceed as before.
@@ IV Resource Estimation
For our encoding, we select the quantum codes constructed from the hypergraph product of classical codes based on (3, 4)-regular bipartite graphs with strong expansion properties Xu et al. (2024). These graphs are selected randomly, and the resulting quantum codes exhibit a minimum encoding rate of 1/25.
Rearrangement of atoms: For a $\big{[}[n,k,d]\big{]}$ hypergraph product (HGP) code where $n$ is the number of physical qubits, $k$ is the number of logical qubits, and $d$ is the code distance, each syndrome extraction (SE) cycle takes $8t_{r}$ seconds Xu et al. (2024), where the factor 8 represents the number of atomic rearrangements needed to complete one error correction cycle, and $t_{r}$ is the time required to move the atomic qubits around for one rearrangement Xu et al. (2024). We express $t_{r}$ as a function of $n$, given by:
@@
where $\tau_{t}=50\mu s$, $a_{p}=0.02\mu m\mu s^{-2}$, $d_{p}=5\mu m$, and $L\approx\sqrt{n}$. The gate times of neutral atom architectures are in the order of microseconds Wintersperger et al. (2023) while $t_{r}$ is in the order of several milliseconds. Therefore $t_{r}$ becomes the dominant rate-limiting time scale and the gate times can be ignored while computing neutral atom qLDPC cycle times.
Logical Failure Rates (LFRs):The Logical Failure Rate (LFR) $R_{L}$ per error correction cycle for a [$n$, $k=n/25$, $d=\sqrt{k}$] HGP code is given by Xu et al. (2024):
@@
where $p_{g}$ is the physical gate error and $n$ is the number of physical qubits, $k=n/25$ is the number of logical qubits, and $d=\sqrt{k}$ is the code distance. RefXu et al. (2024) considers HGP codes such that $k=n/25$ and $d=\sqrt{k}=\sqrt{n}/5$. We consider this configuration as well in our analysis. Similarly, the Logical Failure Rate (LFR) $R_{S}$ per error correction cycle for a surface code of distance $d_{sc}$ is given by Devitt et al. (2016):
@@
Let $n_{a}$ be the number of physical qubits in the ancilla HGP code. Let $n_{t}$ be the number of physical qubits in the qLDPC memory to be transported. Let $n_{c}$ be the number of physical qubits in the qLDPC memory of Charlie Inc.. We make two key assumptions in our analysis. First, we consider the number of physical qubits in Charlie Inc.’s qLDPC memory ($n_{c}$) and the transported qLDPC memory ($n_{t}$) to be equal, denoting both as $n_{m}$, where ’m’ stands for memory. Secondly, we consider the distance of the computational surface code ($d_{sc}$) as a function of the qLDPC memory size, specifically $d_{sc}=\sqrt{n_{m}}/5$, as used by the analysis in Xu et al. (2024).
@@
Bell Pair Creation: Next step, we need to create a Bell-pair. Now, For creating a single Bell Pair, we need two surface codes and $2\frac{\sqrt{n_{m}}}{5}$ error-correction cycles. The Logical Failure Rate (LFR) $R_{1}$ for this is given by Leone et al. (2023); Devitt et al. (2016):
@@
Now for Bell-Measurement, we require $4\frac{\sqrt{n_{m}}}{5}$ QEC cycles Horsman et al. (2012). The LFR $R_{BM}$ for this is given by:
@@
And the fidelity $F_{final}$ of this correlation is given by:
@@
Let $r_{e}$ be the desired target E-Bit rate. Then, the number $N$ of parallel qLDPC blocks required to achieve $r_{e}$ is given by $N=r_{e}/r_{L}$. The blocks get depleted every $T_{4}$ units of time. Therefore, we need to replace $N$ blocks every $T_{4}$ units of time. Let’s say a truck or a train car has the carrying capacity of $B$ blocks, where $B=n_{t}/n_{m}$ where $n_{t}$ is the total number of physical qubits the truck or train car can carry. We need $N/B$ trucks to be released by Charlie Inc. every $T_{4}$ units of time, for one qATM destination. if we are assuming $S$ number of qATM destinations, then the number of trucks $n_{truck}$ released by Charlie Inc. every $T_{4}$ units of time is given by $n_{truck}=SN/B$.

### v2

@@ Abstract
Quantum Entanglement is a vital phenomenon required for realizing secure quantum networks, so much that distributed entanglement can be re-imagined as a commodity which can be traded to enable and maintain these networks. We explore the idea of commercializing entanglement-based cryptography and future applications where advanced quantum memory systems support less advanced users. We design a sneakernet-based quantum communication network with a central party connecting the users through delayed-choice quantum entanglement swapping, using quantum Low-Density-Parity-Check (qLDPC) encoded qubits on neutral atoms. Our analysis compares this approach with traditional surface codes, demonstrating that qLDPC codes offer superior scaling in terms of resource efficiency and logical qubit count. We show that with near-term attainable patch sizes, one can attain medium-to-high fidelity correlations, motivating further research towards the long-term realization of large-scale commercial quantum networks.
@@ I Introduction
The second Quantum Revolution refers to a future technological period when all the features of quantum mechanics are harnessed for information processing devices, expanding beyond the limited application of quantum mechanics in current computers, communication systems, and sensors [1, 2, 3, 4, 5]. The ultimate achievement of this revolution is embodied by quantum computers that are capable of operating on a vast scale, with the ability to tolerate faults and fix errors. These computers would need millions or more logical qubits [6] to tackle complex cryptographic tasks. While the creation of such devices remains in its early stages, there are compelling grounds to anticipate that scalable quantum computers will become tangible at some point within this century [7].
@@
While there is reason to be hopeful about the progress of large-scale quantum computers, it remains uncertain if quantum computers for everyday consumers, such as a quantum smartphone, will ever become feasible. This refers to a compact and powerful computational device that is connected to a quantum network, possibly on a global scale. Possible constraints that could prevent the development of quantum smartphones include the requirement for extremely low temperatures to cool down qubits [8], or the necessity for optical communication channels [9] instead of microwave ones to connect to the quantum internet, the vulnerability of qubits to mechanical vibrations and the impact of the Earth’s magnetic field [10]. Suppose we accept the possibility of large-scale quantum computers and a quantum internet. In that case, it is logical to enquire how a customer, who has no access to a quantum memory, but only has access to conventional gear such as classical detectors, classical memory, and classical computing, can participate in and perform complex quantum protocols.
@@
One way is to do this by the use of distributed quantum entanglement [11]. Quantum entanglement is a pivotal phenomenon that forms the basis of various protocols in quantum communications and quantum computing [12, 13, 14, 15, 16, 17]. Due to it being a major primitive for most quantum tasks, it would be natural to consider quantum entanglement as a tradeable commodity, such that users can purchase it to establish connections with other users in the network to perform quantum tasks such as quantum key distribution, distributed quantum computing, quantum authentication, etc. Commoditizing entanglement is an essential step toward large-scale commercial quantum networks.
@@
To better demonstrate the above-mentioned ideas of entanglement commercialization and quantum access to common users with classical hardware, we design and analyse an entanglement one-time-pad QKD [18] based quantum network model networked using delayed-choice entanglement swapping [19]. In this study, we demonstrate that delayed-choice entanglement swapping enables customers, who only own classical hardware, to use the offerings of a hypothetical future quantum memory company for the purpose of generating secure quantum keys. These keys are based on the E91 protocol [20]. Our design employs a central party Charlie, who generates Bell-pairs, loads the halves separately onto different qLDPC codes, and distributes one set of the halves via sneakernet [21, 22] to multiple endpoints where it is immediately measured upon arrival while keeping the other set of halves to himself on which he later performs Bell measurements to generate correlations between the endpoint measurements of any pair of users. This enables the users to execute this task in a versatile manner by pre-purchasing their key bits prior to selecting their communication partners. This protocol is beneficial for creating secure connections between networks of devices that have minimal hardware needs but can interact with a single, advanced quantum node or several nodes.
@@
We design our protocol on quantum Hyper Graph Product (HGP) codes [23], which are a class of Low-Density Parity-Check (qLDPC) codes [24] for quantum error correction [25], implemented on a neutral atom architecture [26, 27]. These codes offer significant advantages over traditional surface codes [28] in terms of encoding rate and resource efficiency [29, 27]. Our analysis provides a detailed comparison between qLDPC and surface codes, demonstrating that qLDPC codes allow us to achieve higher logical qubit counts with fewer physical qubits as they have constant encoding rates regardless of patch size. This property potentially reduces the hardware requirements and improves the scalability of our proposed system. We have selected neutral atoms as our physical platform, as their ability to shuttle and rearrange within the lattice enables selective qubit unloading and efficient syndrome extraction [27]. This capability also allows us to use only one set of ancilla qubits and a single surface code patch, positioned adjacent to the qLDPC code block, for both loading and unloading operations via teleportation [27].
@@
The foundation of our study is based on the research conducted on consumer-level quantum key distribution [30, 31, 32], as well as the research on quantum sneakernet [21], which explores the spread of entanglement across long distances with high bandwidth and low latency. While our primary focus is on secure communication protocols, it is feasible to adapt our methodology for many applications such as quantum voting [33], clock synchronization [34], and reference frame alignment [35].
@@ II Delayed Choice Entanglement Swapping
To describe our protocol, we start by providing a concise overview of the initial delayed-choice entanglement-swapping technique, as explained by Peres [19]. In accordance with Peres, we examine the most basic iteration of this protocol. This protocol aims to establish entanglement and eventually generate correlated bits between two parties, Alice and Bob, who choose a third party, Charlie, to mediate the entanglement generation. Consider three individuals, namely Alice, Bob, and Charlie, as seen in the circuit diagram in Figure 1.
[…]
Figure 1: Circuit diagram of quantum entanglement swapping. In (a), the conventional entanglement swapping [36] occurs, with Charlie performing the Bell measurement before Alice and Bob does. In (b), the order of measurement is reversed, with Alice and Bob measuring their qubits before Charlie. In both the circuits, all the qubits are starting at |$\ket{0}$ state. These two circuits are equivalent as the two sets of measurements are parallel, non-overlapping, and independent of each other.
Two Bell Pairs $q_{0}q_{1}$ and $q_{2}q_{3}$ are present such that $q_{0}$ is in the custody of Alice, $q_{3}$ with Bob, and $q_{1}$ and $q_{2}$ with Charlie. Qubits $q_{0}$ of Alice and $q_{3}$ of Bob get entangled when Charlie performs a joint Bell-measurement on his qubits $q_{1}$ and $q_{2}$. However, the specific relationship between Alice and Bob’s qubits cannot be known until Charlie reveals the outcome of his Bell state measurement. Alice and Bob’s photons are perfectly entangled, disregarding any errors or decoherence. This exact entanglement allows them to be used for various entanglement-based protocols, such as cryptography using the E91 protocol [20].
@@
Peres’ idea goes beyond entanglement swapping by recognizing that Alice, Bob, and Charlie might potentially be space-like separated. This suggests that one can discover a frame of reference when the sequence of qubit measurement is modified, as shown by the Gisin group [37]. In the preceding discourse, we used the assumption that Charlie conducted his measurement before Alice and Bob did. Alternatively, it is possible for Alice and Bob to conduct their measurements prior to Charlie (Figure 1(b)). In this scenario, the outcome of their measurements itself is not affected by whether or not Charlie performed either the Bell state or independent measurement, but the degree of correlation between Alice and Bob’s measured bits is dependent upon the kind of measurement performed by Charlie. This serves as the foundation for the delayed-choice entanglement-swapping protocol and is crucial for the advancement of our ideas. In the delayed-choice entanglement swapping technique, Alice and Bob perform measurements on their qubits and record the outcomes using classical storage. At a later point, Charlie has a decision to either measure his qubits separately or do a Bell state measurement. The measurement outputs are either associated in a known manner or wholly uncorrelated, depending on the measurement option chosen by Charlie and the random basis choice of Alice and Bob.
@@ III Setup
Figure 2: (Above) A schematic of a 5-node delayed choice quantum network with Charlie Inc. in the center. (Below) Delayed choice entanglement swapping protocol between two nodes Alice and Bob. S represents the surface code patches, A refers to the ancilla patches encoded in HGP code, C denotes Charlie’s qLDPC memory and T represents the qLDPC memory transported to the end-nodes. Here, E represents the initial entangled Bell-pair and B denotes bell measurement.
@@
and is also described concisely in pseudocode 1. Charlie Inc., acting as a central node, first prepares Bell pairs in surface codes and teleports each half into separate qLDPC code blocks via a qLDPC (HGP) ancilla. This is done because direct computations and measurements for qLDPC codes are not yet well defined at the time of this writing. Charlie Inc. then keeps one set of neutral-atom encoded qLDPC blocks and sends the other set to an end-user via physical transport, like a train or a truck. This process is done for several such end-users at different destinations. This is called the ”sneakernet” method. At the user’s end, the user purchases a certain number of bits based on their needs. At the user’s end, a quantum ATM (qATM) “unloads” the incoming qLDPC blocks by sequentially teleporting the constituent qubits onto a surface code, followed by measuring the surface code randomly in X or Z bases, and commercially selling the resulting classical bits to users, which is in turn stored in a device belonging to the user, such as a smartphone. For our analysis, we assume that the process of transferring classical bits onto the end-users’ classical devices is secure. The measurement bases are then made public.
@@
Among several end-users in the network, let’s say two of them, namely Alice and Bob, want to connect at a later time. They decide upon how many bits they need to correlate for their requirement. Charlie Inc. then converts the corresponding logical qubits in the qLDPC blocks on his end into surface code patches and performs Bell measurements on those patches, thereby correlating Alice and Bob’s bits. Finally, Alice and Bob use their correlated bits to generate a draft encryption key. Depending on their security requirement, they can enhance the key’s security using CHSH tests [38] and privacy amplification, sacrificing some bits in the process. The remaining bits form a secure encryption key for end-to-end message encryption. This approach turns entanglement into a tradeable commodity and allows users with only classical hardware to participate in quantum cryptography protocols. For example, Alice measures a number of qubits, labeled ${A_{1},A_{2},...,A_{n}}$ and informs Charlie Inc. which qubit measurement results she now owns. Charlie Inc. then records the owner of the qubits and stores the other half of the Bell pairs in long-term quantum storage. Similarly, Bob buys the rights to measure qubits ${B_{1},B_{2},...,B_{n}}$ and Dave buys the rights to measure qubits ${D1,D_{2},...,D_{n}}$. At this stage, neither participant has decided who they wish to exchange correlations with, they have simply ‘topped up’ their future quantum potential i.e. the bits with which they can perform an unconditionally secure one-time-pad with another user of their choice in the future. If Alice and Dave wish to generate entangled data for some protocol, then they need to publicly declare to Charlie Inc. how many topped up bits they want to have correlated, then Charlie Inc. performs the entanglement swap between the appropriate (stored) Bell pairs and broadcasts the measurement results. If Alice and Dave wish to check Charlie Inc.’s fidelity, they may sacrifice some of their bits, and assuming they are satisfied, can proceed with standard key distillation and privacy amplification [39]. If Alice then wishes to share secrets with Bob, then again, she need only declare to Charlie Inc. how many bits she wants to correlate (of course, she cannot use measurement results that have previously been used with Dave), and proceed as before.
@@ IV Resource Estimation
For our protocol, we select the quantum LDPC codes constructed from the hypergraph product of classical codes based on (3, 4)-regular bipartite graphs with strong expansion properties [27]. These graphs are selected randomly, and the resulting quantum codes exhibit a minimum encoding rate of 1/25. In ref. [27], the authors discuss the implementation of computation with qLDPC codes using neutral atoms, which provide an order of magnitude improvement in the number of physical qubits used compared to surface codes. They introduce a technique by which logical qubits can be teleported between the surface codes and qLDPC code blocks, thus aiding in the possibility of computation and individual logical qubit measurements, which becomes essential for the implementation for our protocol. The authors also provide closed-form expressions for various parameters such as logical failure rates and cycle times, which we incorporate in our resource analysis.
Rearrangement of atoms: For a $\big{[}[n,k,d]\big{]}$ hypergraph product (HGP) code with a coding rate $r=k/n$ where $n$ is the number of physical qubits, $k$ is the number of logical qubits, and $d$ is the code distance, each syndrome extraction (SE) cycle takes time $8t_{r}$ [27], where the factor 8 represents the number of atomic rearrangements needed to complete one error correction cycle, and $t_{r}$ is the time required to move the atomic qubits around for one rearrangement [27]. By overlapping the syndrome measurements with the atom repositioning of the next iteration, we can avoid extending the total execution time. We express $t_{r}$ as a function of $n$, given by:
@@
where $\tau_{t}=50\mu s$ is the transfer time between the atomic traps, $a_{p}=0.02\mu m\mu s^{-2}$ is the maximum acceleration rate of the atom, $d_{p}=5\mu m$ is the spacing between the atoms in the qLDPC grid, and $L\approx\sqrt{n}$ is the number of atoms in a line to be rearranged. The gate times of neutral atom architectures are in the order of microseconds [40] while $t_{r}$ is in the order of several milliseconds. Therefore $t_{r}$ becomes the dominant rate-limiting time scale and the gate times can be ignored while computing neutral atom qLDPC cycle times.
Logical Failure Rates (LFRs):The Logical Failure Rate (LFR) $R_{L}$ per error correction cycle for a [$n$, $k=n/25$, $d=\sqrt{k}$] HGP code is given by [27]:
@@
where $p_{g}$ is the physical gate error and $n$ is the number of physical qubits, $k=n/25$ is the number of logical qubits, and $d=\sqrt{k}$ is the code distance. Ref[27] considers HGP codes such that $k=n/25$ and $d=\sqrt{k}=\sqrt{n}/5$. We consider this configuration as well in our analysis. Similarly, the Logical Failure Rate (LFR) $R_{S}$ per error correction cycle for a surface code of distance $d_{sc}$ is given by [21]:
@@
Let $n_{a}$ be the number of physical qubits in the ancilla HGP code. Let $n_{t}$ be the number of physical qubits in the qLDPC memory to be transported. Let $n_{c}$ be the number of physical qubits in the qLDPC memory of Charlie Inc.. We make two key assumptions in our analysis. First, we consider the number of physical qubits in Charlie Inc.’s qLDPC memory ($n_{c}$) and the transported qLDPC memory ($n_{t}$) to be equal, denoting both as $n_{m}$, where ’m’ stands for memory. Secondly, we consider the distance of the computational surface code ($d_{sc}$) as a function of the qLDPC memory size, specifically $d_{sc}=\sqrt{n_{m}}/5$, as used by the analysis in [27].
@@
Bell Pair Creation: Next step, we need to create a Bell-pair. Now, For creating a single Bell Pair, we need two surface codes and $2\frac{\sqrt{n_{m}}}{5}$ error-correction cycles. The Logical Failure Rate (LFR) $R_{1}$ for this is given by [41, 21]:
@@
Now for Bell-Measurement, we require $4\frac{\sqrt{n_{m}}}{5}$ QEC cycles [42]. The LFR $R_{BM}$ for this is given by:
@@
And the success probability $F_{final}$ of this correlation is given by:
@@
Let $r_{e}$ be the desired target entangled bit (E-Bit) rate. Then, the number $N$ of parallel qLDPC blocks required to achieve $r_{e}$ is given by $N=r_{e}/r_{L}$. The blocks get depleted every $T_{4}$ units of time. Therefore, we need to replace $N$ blocks every $T_{4}$ units of time. Let’s say a truck or a train car has the carrying capacity of $B$ blocks, where $B=n_{t}/n_{m}$ where $n_{t}$ is the total number of physical qubits the truck or train car can carry. We need $N/B$ trucks to be released by Charlie Inc. every $T_{4}$ units of time, for one qATM destination. if we are assuming $S$ number of qATM destinations, then the number of trucks $n_{truck}$ released by Charlie Inc. every $T_{4}$ units of time is given by $n_{truck}=SN/B$.